OpenStudy (anonymous):

What is the rule for this function: range {-3,-2,-1,0,1} domain {5,2,-1,-4,-7}

OpenStudy (anonymous):

3x-4

OpenStudy (anonymous):

sorry, -3x-4

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

no prob. in situations like these it is often helpful to look at the zero case, to see what might be added or subtracted

jimthompson5910 (jim_thompson5910):

As x decreases by 3 each time, y increases by 1 each time. So the slope is m = 1/(-3) = -1/3 Now pick any ordered pair, say (5,-3). So x = 5 and y=-3 Plug these values into y-y1=m(x-x1) y-(-3)=(-1/3)(x-5) y+3=(-1/3)(x-5) y+3=(-1/3)x-(1/3)(-5) y+3=(-1/3)x+5/3 y=(-1/3)x+5/3-3 y=(-1/3)x-4/3 So the equation is \[\large y=-\frac{1}{3}x-\frac{4}{3}\]

OpenStudy (anonymous):

Wow--Thank you!